OpenStudy (anonymous):

Solve the following equation, answer as a reduced, mixed number. 6(x - 1) = 4(x - 2)

OpenStudy (aperez):

Algebra?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

6x-6= 4x-8 6x-4x=6-8 2x=-2 x=-2/2 x= -1

OpenStudy (anonymous):

Thank you, roselin. At least some people on here are helpful...

OpenStudy (anonymous):

If you want the answer, not problem-solving skills.